What does this reaction produce?
B4C + CO2—>
1
Expert's answer
2018-10-26T08:20:09-0400

Solution

B4C doesn't react with CO2.

It reacts with O2, H2O

B4C+4O2= 2B2O3+ CO2

B4C+ H2O= CO + B2O3